6. Find the ratio of milk with 3.8% fat to milk with 0.5% fat that have to be mixed in order to produce a blend with 3.5% fat.
Answer #1

Let 3.8% fat milk added be equal to x.

So, 0.5% fat milk added will be (1-x).

3.8x + 0.5(1-x) = 3.5

3.8x + 0.5 - 0.5x = 3.5

3.3x = 3

x = 3 / 3.3

= 0.909

1 - x = 0.091

Ratio = 0.909 : 0.091 = 90.9 : 9.1 (in %)